# refrences.bib

### v1

@@ 5 Local and global optimal solutions
As mentioned before, throughout this section we assume $f$ is a continuous function, that is nondecreasing (non-increasing) in $x_{j}$, $\forall j\in\mathscr{J}^{+}$ ($\forall j\in\mathscr{J}^{-}$).
Definition 8.
@@
    Let $x\in S\left(A^{+},A^{-},b\right)$ be an arbitrary feasible solution. We shall show that $f\left(x^{*}\left(e^{*}\right)\right)\leq f(x)$. From Theorem 1, $x\in S\left(e_{0}\right)$ for some $e_{0}\in E$. So, theorem 2 requires $f\left(x^{*}\left(e_{0}\right)\right)\leq f(x)$, which together with $f\left(x^{*}\left(e^{*}\right)\right)\leq f\left(x^{*}\left(e_{0}\right)\right)$ imply $f\left(x^{*}\left(e^{*}\right)\right)\leq f(x)$. ∎
From Theorem 2, $x^{*}(e)$ is a local optimal solution, $\forall e\in E$. So, Theorem 3 provides also a necessary optimality condition in the sense that if $x^{*}$ is an optimal solution of Problem (1), then it must belong to $F^{*}$. In other words, each solution $x^{*}(e)\in F^{*}$ is an optimal candidate solution. Additionally, the global optimal value is obtained as $f\left(x^{*}\left(e^{*}\right)\right)=\min\left\{x^{*}(e):e\in E\right\}$.
Example 4 (Linear functions).
Assume that the objective function of Problem (1) is defined as the linear form $f(x)=\sum_{j=1}^{n}c_{j}x_{j}$, where $c_{j}\in\mathbb{R}$, $\forall j\in\mathscr{J}$. So, it is clear that $\mathscr{J}^{+}=\left\{j\in\mathscr{J}:c_{j}\geq 0\right\}$ and $\mathscr{J}^{-}=\left\{j\in\mathscr{J}:c_{j}<0\right\}$. For instance, consider the problem stated in Examples $1-3$ with the following objective function:
@@
In this example, we have $\mathscr{J}^{+}=\{1,2,5,6,8\}$ and $\mathscr{J}^{-}=\{3,4,7,9\}$. To calculate $x^{*}\left(e_{1}\right)$, we note that $e_{1}=[6,6]$, and therefore $e_{1}(1)=6$ and $e_{1}(2)=6$ (see Table 5). However, by noting that the sixth column in Table 5 corresponds to the eighth column of the main problem, it follows that $\mathscr{I}_{8}\left(e_{1}\right)=\{1,2\}$. Moreover, $\mathscr{I}_{j}\left(e_{1}\right)=\varnothing$, $\forall j\in\{1,\ldots,7\}\bigcup\{9\}$. Hence, from Definition 8 (relation (9)), it is concluded that
@@
Also, the objective values of the above local optimal solutions are computed as $f_{1}\left(x^{*}\left(e_{1}\right)\right)=-0.9$, $f_{1}\left(x^{*}\left(e_{2}\right)\right)=-3.6$, $f_{1}\left(x^{*}\left(e_{3}\right)\right)=-1.3$ and $f_{1}\left(x^{*}\left(e_{4}\right)\right)=-3.3$. So, by Theorem 3, the global minimum solution of the problem is $x^{*}\left(e_{2}\right)$ with an optimal value of $f_{1}\left(x^{*}\left(e_{2}\right)\right)=-3.6$.
Example 5 (Support function of a set).
@@
with domain dom $\text{{dom}~{}}f(x)=\left\{x\in\mathbb{R}^{n}:\sup\left\{x^{T}y:y\in\mathbb{C}\right\}<\infty\right\}$, where $\mathbb{C}\subseteq\mathbb{R}^{n}$ and $\mathbb{C}\neq\varnothing$. Many theoretical and applied aspects of this function can be found in [1, 2]. If $x$ is an arbitrary fixed point in $\mathbb{R}^{n}$ and $\mathbb{C}$ is a non-empty bounded polyhedral set (polytope), then $\sup\left\{x^{T}y:y\in\mathbb{C}\right\}$ becomes equivalent to the following traditional linear programming problem:
@@
where $x_{j}$’s and $y_{j}$’s ($j\in\mathscr{J}$) are the cost coefficients and the decision variables, respectively. Now, suppose that the cost vector $x$ is perturbed along the $j$’th unit vector $u_{j}$ (a vector having zero components, except for a 1 in the $j$’th position), that is, $x$ is replaced by $x+\lambda u_{j}$ where $\lambda\geq 0$ In other words, we increase $x_{j}$ to $x_{j}+\lambda$ and hold other cost coefficients fixed. However, since the parametric analysis on the cost vector in a linear programming problem always produces a piecewise-linear and increasing function $z(\lambda)$ [1], it is concluded that $f_{2}$ is non-decreasing in variable $x_{j}$, $\forall j\in\mathscr{J}$. Consequently, we have $\mathscr{J}^{+}=\mathscr{J}$ and $\mathscr{J}^{-}=\varnothing$. Therefore, by considering the problem stated in Examples $1-3$, the local optimal solutions $x^{*}\left(e_{1}\right)$, $x^{*}\left(e_{2}\right)$, $x^{*}\left(e_{3}\right)$ and $x^{*}\left(e_{4}\right)$ are obtained by (9) as follows:
@@
So, $f_{2}\left(x^{*}\left(e_{1}\right)\right)=0.8$ and in the same way, we get $f_{2}\left(x^{*}\left(e_{2}\right)\right)=0.75$, $f_{2}\left(x^{*}\left(e_{3}\right)\right)=0.75$ and $f_{2}\left(x^{*}\left(e_{4}\right)\right)=0.75$ Hence, $x^{*}(e_{2})$, $x^{*}(e_{3})$ and $x^{*}(e_{4})$ are the global optimal solutions with the optimal value of 0.75.
Example 6 (perspective function).
@@
It is to be noted that since $I_{9}=[0.2,1]$ (from Table 3), then $0<0.2\leq x_{9}$ for any $x\in S\left(A^{+},A^{-},b\right)$. So, $f_{3}$ is indeed the perspective function of $\left|x_{1}\right|^{p}+\cdots+\left|x_{8}\right|^{p},p\geq 1$ [2]. In this case, $\mathscr{J}^{+}=\{1,\ldots,8\}$ and $\mathscr{J}^{-}=\{9\}$. Based on Definition 8 , the local optimal solutions $x^{*}\left(e_{1}\right)$, $x^{*}\left(e_{2}\right)$, $x^{*}\left(e_{3}\right)$ and $x^{*}\left(e_{4}\right)$ are obtained for $p=3$ as follows:
@@ Conclusion
In this paper, the bipolar fuzzy relational equations were developed to the most general case in which the fuzzy compositions are defined by an arbitrary continuous t-norm. An algorithm was proposed for finding a global optimal solution to a wide class of nonlinear optimization problems constrained by the continuous bipolar fuzzy relation equations. The analytical properties of the continuous bipolar FREs were investigated and the feasible solutions set of the problem was completely characterized by a finite number of compact sets that may not be connected sets. Two necessary feasibility conditions and a necessary and sufficient condition were also presented to determine the feasibility of the problem. In addition, five simplification rules were proposed to speed up solution finding.
@@
{tblr}
c—c—c t-norm & Function Parameter
Minimum $T_{M}(x,y)=\min\{x,y\}$
Product $T_{P}(x,y)=xy$
Einstein Product $T_{EP}(x,y)=\frac{xy}{2-(x+y-xy)}$
Lukasiewicz $T_{L}(x,y)=max\{0,x+y-1\}$
@@
Table A2: Values $l$ and $u$ discussed in Lemma 1 for the t-norms stated in Table A1. The values $l$ and $u$ are presented for the cases that $a\geq b>0$ and $a>b\geq 0$, respectively; because, for all the cases, it is easily verified that $l=0$ if $b=0$, and $u=1$ if $a=b$.
{tblr}
c—c—c t-norm $l(a\geq b>0)$ $u(a>b\geq 0)$
@@
    From Theorem 1, $S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)=\bigcup_{e^{\prime}\in E^{\prime}}S\left(e^{\prime}\right)$, where $E^{\prime}$ is the set of all the restrictions of admissible functions $e\in E$ to $\mathscr{I}-\left\{i_{0}\right\}$. So, for each $x\in S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)$, there exists at least one $e_{0}^{\prime}\in E^{\prime}$ such that $x\in S\left(e_{0}^{\prime}\right)$. Now, from (5), for each $j\in\mathscr{J}$ we have either $x_{j}\in\bigcap_{i\in\mathscr{I}_{j}\left(e_{0}^{\prime}\right)}S_{ij}^{\prime}$ or $x_{j}\in I_{j}$. But, in the former case, since $S_{ij}^{\prime}=S_{ij}\cap I_{j}$ (Definition 4), we have again $x_{j}\in\bigcap_{i\in\mathscr{I}_{j}\left(e_{0}^{\prime}\right)}S_{ij}^{\prime}\subseteq I_{j}$. Consequently, $x_{j}\in I_{j}$ (and therefore $x_{j}\in I_{i_{0}j}$ from Definition 4), $\forall j\in\mathscr{J}$. ∎
  • The proof of Rule 1.
@@
    Let $x\in S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)$. We shall show that $x$ also satisfies the $i_{0}$’th equation. From Lemma B1, $x_{j}\in I_{j}$, $\forall j\in\mathscr{J}\left(*_{1}\right)$. Also, since $S\left(A^{+},A^{-},b\right)\neq\varnothing$, there exists at least one $j_{0}\in\mathscr{J}$ such that $S_{i_{0}j_{0}}^{\prime}\neq\varnothing$ (Lemma 3 (b)). On the other hand, since $b_{i_{0}}=0$, then we have $S_{i_{0}j_{0}}^{\prime}=I_{i_{0}j_{0}}$ which together with $I_{j_{0}}\subseteq I_{i_{0}j_{0}}$ (Definition 4) and (*1) imply $x_{j_{0}}\in S_{i_{0}j_{0}}^{\prime}$ (*2). Now, the result follows from (*1),(*2) and Corollary 2. ∎
  • The proof of Rule 2.
@@
    For any $x\in S\left(A^{+},A^{-},b\right)$, Lemma 4 implies $x_{j_{0}}\in I_{j_{0}}=\{k\}$, i.e., $x_{j_{0}}=k$. Now, from Lemma B1, if $x\in S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)$, then $x_{j}\in I_{i_{0}j}$, $\forall j\in\mathscr{J}$(*1). But, since $k\in S_{i_{0}j_{0}}^{\prime}$, then we have $S_{i_{0}j_{0}}^{\prime}\neq\varnothing$, that together with $S_{i_{0}j_{0}}^{\prime}=S_{i_{0}j_{0}}\cap I_{j_{0}}$ and $I_{j_{0}}=\{k\}$ imply $S_{i_{0}j_{0}}^{\prime}=\{k\}$. Consequently, $x_{j_{0}}\in S_{i_{0}j_{0}}^{\prime}$ (*2). Hence, from (*1),(*2) and Corollary 2, it follows that $x$ satisfies the $i_{0}$’th equation. ∎
  • The proof of Rule 3.
@@
    If $x\in S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)$, then by Lemma B1, we conclude that $x_{j}\in I_{i_{0}j}$, $\forall j\in\mathscr{J}$ (*1). Also, from Lemma 4, there exists at least one $j_{i}\in\mathscr{J}$ such that $x_{j_{i}}\in S_{ij_{i}}^{\prime}$. But, since $S_{ij}^{\prime}\subseteq S_{i_{0}j}^{\prime}$, then we have $x_{j_{i}}\in S_{i_{0}j_{i}}^{\prime}$ (*2). So, (*1), (*2) and Corollary 2 imply that $x$ satisfies the $i_{0}$’th equation. ∎
  • The proof of Rule 4.
@@
    If $x\in S\left(A^{+},A^{-},b\right)$, then from Lemma 4, there exists at least one $j\in\mathscr{J}$ such that $x_{j}\in S_{i_{0}j}^{\prime}$. But, since $\mathscr{J}_{i_{0}}=\left\{j_{0}\right\}$, we have necessarily $x_{j_{0}}\in S_{i_{0}j_{0}}^{\prime}=\{k\}$, that means $x_{j_{0}}=k$. Moreover, suppose that $i\in\mathscr{I}$ and $k\in S_{ij_{0}}^{\prime}$. If $x\in S_{\{i\}}\left(A^{+},A^{-},b\right)$, then $x$ satisfies the $i_{0}$’th equation, and therefore we must have $x_{j_{0}}=k\in S_{ij_{0}}^{\prime}$. Also, from Lemma B1, $x_{j}\in I_{ij}$, $\forall j\in\mathscr{J}$. Hence, Corollary 2 implies that $x$ also satisfies the $i$’th equation. ∎
  • The proof of Rule 5.
@@
    Let $x\in S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)$. From Lemma B1, $x_{j_{0}}\in I_{j_{0}}$ and $x_{j}\in I_{i_{0}j}$, $\forall j\in\mathscr{J}$. But, since $S_{i_{0}j_{0}}^{\prime}=I_{j_{0}}$, then $x_{j_{0}}\in I_{j_{0}}=S_{i_{0}j_{0}}^{\prime}$, and therefore from Corollary 2 it follows that $x$ satisfies the $i_{0}$’th equation. ∎
[…]

### v2

@@ 5 Local and global optimal solutions
As mentioned before, throughout this section we assume $f$ is a continuous function, that is non-decreasing (non-increasing) in $x_{j}$, $\forall j\in\mathscr{J}^{+}$ ($\forall j\in\mathscr{J}^{-}$).
Definition 5.1.
@@
Let $x\in S\left(A^{+},A^{-},b\right)$ be an arbitrary feasible solution. We shall show that $f\left(x^{*}\left(e^{*}\right)\right)\leq f(x)$. From Theorem 3.9, $x\in S\left(e_{0}\right)$ for some $e_{0}\in E$. So, theorem 5.2 requires $f\left(x^{*}\left(e_{0}\right)\right)\leq f(x)$, which together with $f\left(x^{*}\left(e^{*}\right)\right)\leq f\left(x^{*}\left(e_{0}\right)\right)$ imply $f\left(x^{*}\left(e^{*}\right)\right)\leq f(x)$. ∎
From Theorem 5.2, $x^{*}(e)$ is a local optimal solution, $\forall e\in E$. So, Theorem 5.3 provides also a necessary optimality condition in the sense that if $x^{*}$ is an optimal solution of Problem (1), then it must belong to $F^{*}$. In other words, each solution $x^{*}(e)\in F^{*}$ is an optimal candidate solution. Additionally, the global optimal value is obtained as $f\left(x^{*}\left(e^{*}\right)\right)=\min\left\{x^{*}(e):e\in E\right\}$.
Example 5.4 (Linear functions).
Assume that the objective function of Problem (1) is defined as the linear form $f(x)=\sum_{j=1}^{n}c_{j}x_{j}$, where $c_{j}\in\mathbb{R}$, $\forall j\in\mathscr{J}$. So, it is clear that $\mathscr{J}^{+}=\left\{j\in\mathscr{J}:c_{j}\geq 0\right\}$ and $\mathscr{J}^{-}=\left\{j\in\mathscr{J}:c_{j}<0\right\}$. For instance, consider the problem stated in Examples 2.9,3.10 and 4.6 with the following objective function:
@@
In this example, we have $\mathscr{J}^{+}=\{1,2,5,6,8\}$ and $\mathscr{J}^{-}=\{3,4,7,9\}$. To calculate $x^{*}\left(e_{1}\right)$, we note that $e_{1}=[6,6]$, and therefore $e_{1}(1)=6$ and $e_{1}(2)=6$ (see Table 5). However, by noting that the sixth column in Table 5 corresponds to the eighth column of the main problem, it follows that $\mathscr{I}_{8}\left(e_{1}\right)=\{1,2\}$. Moreover, $\mathscr{I}_{j}\left(e_{1}\right)=\varnothing$, $\forall j\in\{1,\ldots,7\}\bigcup\{9\}$. Hence, from Definition 5.1 (relation (9)), it is concluded that
@@
Also, the objective values of the above local optimal solutions are computed as $f_{1}\left(x^{*}\left(e_{1}\right)\right)=-0.9$, $f_{1}\left(x^{*}\left(e_{2}\right)\right)=-3.6$, $f_{1}\left(x^{*}\left(e_{3}\right)\right)=-1.3$ and $f_{1}\left(x^{*}\left(e_{4}\right)\right)=-3.3$. So, by Theorem 5.3, the global minimum solution of the problem is $x^{*}\left(e_{2}\right)$ with an optimal value of $f_{1}\left(x^{*}\left(e_{2}\right)\right)=-3.6$.
Example 5.5 (Support function of a set).
@@
with domain $\text{{dom}~}f(x)=\left\{x\in\mathbb{R}^{n}:\sup\left\{x^{T}y:y\in\mathbb{C}\right\}<\infty\right\}$, where $\mathbb{C}\subseteq\mathbb{R}^{n}$ and $\mathbb{C}\neq\varnothing$. Many theoretical and applied aspects of this function can be found in [1, 2]. If $x$ is an arbitrary fixed point in $\mathbb{R}^{n}$ and $\mathbb{C}$ is a non-empty bounded polyhedral set (polytope), then $\sup\left\{x^{T}y:y\in\mathbb{C}\right\}$ becomes equivalent to the following traditional linear programming problem:
@@
where $x_{j}$’s and $y_{j}$’s ($j\in\mathscr{J}$) are the cost coefficients and the decision variables, respectively. Now, suppose that the cost vector $x$ is perturbed along the $j$’th unit vector $u_{j}$ (a vector having zero components, except for a 1 in the $j$’th position), that is, $x$ is replaced by $x+\lambda u_{j}$ where $\lambda\geq 0$ . In other words, we increase $x_{j}$ to $x_{j}+\lambda$ and hold other cost coefficients fixed. However, since the parametric analysis on the cost vector in a linear programming problem always produces a piecewise-linear and increasing function $z(\lambda)$ [1], it is concluded that $f_{2}$ is non-decreasing in variable $x_{j}$, $\forall j\in\mathscr{J}$. Consequently, we have $\mathscr{J}^{+}=\mathscr{J}$ and $\mathscr{J}^{-}=\varnothing$. Therefore, by considering the problem stated in Examples 2.9,3.10 and 4.6, the local optimal solutions $x^{*}\left(e_{1}\right)$, $x^{*}\left(e_{2}\right)$, $x^{*}\left(e_{3}\right)$ and $x^{*}\left(e_{4}\right)$ are obtained by (9) as follows:
@@
So, $f_{2}\left(x^{*}\left(e_{1}\right)\right)=0.8$ and in the same way, we get $f_{2}\left(x^{*}\left(e_{2}\right)\right)=0.75$, $f_{2}\left(x^{*}\left(e_{3}\right)\right)=0.75$ and $f_{2}\left(x^{*}\left(e_{4}\right)\right)=0.75$ . Hence, $x^{*}(e_{2})$, $x^{*}(e_{3})$ and $x^{*}(e_{4})$ are the global optimal solutions with the optimal value of 0.75.
Example 5.6 (perspective function).
@@
It is to be noted that since $I_{9}=[0.2,1]$ (from Table 3), then $0<0.2\leq x_{9}$ for any $x\in S\left(A^{+},A^{-},b\right)$. So, $f_{3}$ is indeed the perspective function of $\left|x_{1}\right|^{p}+\cdots+\left|x_{8}\right|^{p},p\geq 1$ [2]. In this case, $\mathscr{J}^{+}=\{1,\ldots,8\}$ and $\mathscr{J}^{-}=\{9\}$. Based on Definition 5.1, the local optimal solutions $x^{*}\left(e_{1}\right)$, $x^{*}\left(e_{2}\right)$, $x^{*}\left(e_{3}\right)$ and $x^{*}\left(e_{4}\right)$ are obtained for $p=3$ as follows:
@@ Conclusion
This paper develops bipolar fuzzy relational equations for the most general case in which fuzzy compositions are defined by arbitrary continuous t-norms. An algorithm was proposed for finding a global optimal solution to a wide class of nonlinear optimization problems constrained by continuous bipolar fuzzy relation equations. The analytical properties of the continuous bipolar FREs were investigated, and the feasible solution set of the problem was completely characterized by a finite number of compact sets that may not be connected sets. Two necessary feasibility conditions and a necessary and sufficient condition were also presented to determine the feasibility of the problem. In addition, five simplification rules were proposed to speed up solution finding. Furthermore, as mentioned earlier, the proposed method provides some mathematical background that can be used to develop improved methods for solving optimization problems that are more likely to have different types of objective functions rather than decreasing (non-increasing) functions. Despite this, since the complexity of solving the problem increases exponentially with the problem size, it is always a challenge to find other simplification rules or methods that can reach the optimal solution.
@@
t-norm Function Parameter
Minimum $T_{M}(x,y)=\min\{x,y\}$ -
Product $T_{P}(x,y)=xy$ -
Einstein Product $T_{EP}(x,y)=\frac{xy}{2-(x+y-xy)}$ -
Lukasiewicz $T_{L}(x,y)=max\{0,x+y-1\}$ -
@@
Table A2: Values $l$ and $u$ discussed in Lemma 2.2 for the t-norms stated in Table A1. The values $l$ and $u$ are presented for the cases that $a\geq b>0$ and $a>b\geq 0$, respectively; because, for all the cases, it is easily verified that $l=0$ if $b=0$, and $u=1$ if $a=b$.
t-norm $l(a\geq b>0)$ $u(a>b\geq 0)$
@@
From Theorem 3.9, $S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)=\bigcup_{e^{\prime}\in E^{\prime}}S\left(e^{\prime}\right)$, where $E^{\prime}$ is the set of all the restrictions of admissible functions $e\in E$ to $\mathscr{I}-\left\{i_{0}\right\}$. So, for each $x\in S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)$, there exists at least one $e_{0}^{\prime}\in E^{\prime}$ such that $x\in S\left(e_{0}^{\prime}\right)$. Now, from (5), for each $j\in\mathscr{J}$ we have either $x_{j}\in\bigcap_{i\in\mathscr{I}_{j}\left(e_{0}^{\prime}\right)}S_{ij}^{\prime}$ or $x_{j}\in I_{j}$. But, in the former case, since $S_{ij}^{\prime}=S_{ij}\cap I_{j}$ (Definition 2.7), we have again $x_{j}\in\bigcap_{i\in\mathscr{I}_{j}\left(e_{0}^{\prime}\right)}S_{ij}^{\prime}\subseteq I_{j}$. Consequently, $x_{j}\in I_{j}$ (and therefore $x_{j}\in I_{i_{0}j}$ from Definition 2.7), $\forall j\in\mathscr{J}$. ∎
The Proof of Rule 4.1.
@@
Let $x\in S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)$. We shall show that $x$ also satisfies the $i_{0}$’th equation. From Lemma B1, $x_{j}\in I_{j}$, $\forall j\in\mathscr{J}$(*1). Also, since $S\left(A^{+},A^{-},b\right)\neq\varnothing$, there exists at least one $j_{0}\in\mathscr{J}$ such that $S_{i_{0}j_{0}}^{\prime}\neq\varnothing$ (Lemma 3.1 (b)). On the other hand, since $b_{i_{0}}=0$, then we have $S_{i_{0}j_{0}}^{\prime}=I_{i_{0}j_{0}}$ which together with $I_{j_{0}}\subseteq I_{i_{0}j_{0}}$ (Definition 2.7) and (*1) imply $x_{j_{0}}\in S_{i_{0}j_{0}}^{\prime}$ (*2). Now, the result follows from (*1),(*2) and Corollary 3.3. ∎
The Proof of Rule 4.2.
@@
For any $x\in S\left(A^{+},A^{-},b\right)$, Lemma 3.2 implies $x_{j_{0}}\in I_{j_{0}}=\{k\}$, i.e., $x_{j_{0}}=k$. Now, from Lemma B1, if $x\in S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)$, then $x_{j}\in I_{i_{0}j}$, $\forall j\in\mathscr{J}$(*1). But, since $k\in S_{i_{0}j_{0}}^{\prime}$, then we have $S_{i_{0}j_{0}}^{\prime}\neq\varnothing$, that together with $S_{i_{0}j_{0}}^{\prime}=S_{i_{0}j_{0}}\cap I_{j_{0}}$ and $I_{j_{0}}=\{k\}$ imply $S_{i_{0}j_{0}}^{\prime}=\{k\}$. Consequently, $x_{j_{0}}\in S_{i_{0}j_{0}}^{\prime}$ (*2). Hence, from (*1),(*2) and Corollary 3.3, it follows that $x$ satisfies the $i_{0}$’th equation. ∎
The Proof of Rule 4.3.
@@
If $x\in S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)$, then by Lemma B1, we conclude that $x_{j}\in I_{i_{0}j}$, $\forall j\in\mathscr{J}$ (*1). Also, from Lemma 3.2, there exists at least one $j_{i}\in\mathscr{J}$ such that $x_{j_{i}}\in S_{ij_{i}}^{\prime}$. But, since $S_{ij}^{\prime}\subseteq S_{i_{0}j}^{\prime}$, then we have $x_{j_{i}}\in S_{i_{0}j_{i}}^{\prime}$ (*2). So, (*1), (*2) and Corollary 3.3 imply that $x$ satisfies the $i_{0}$’th equation. ∎
The Proof of Rule 4.4.
@@
If $x\in S\left(A^{+},A^{-},b\right)$, then from Lemma 3.2, there exists at least one $j\in\mathscr{J}$ such that $x_{j}\in S_{i_{0}j}^{\prime}$. But, since $\mathscr{J}_{i_{0}}=\left\{j_{0}\right\}$, we have necessarily $x_{j_{0}}\in S_{i_{0}j_{0}}^{\prime}=\{k\}$, that means $x_{j_{0}}=k$. Moreover, suppose that $i\in\mathscr{I}$ and $k\in S_{ij_{0}}^{\prime}$. If $x\in S_{\{i\}}\left(A^{+},A^{-},b\right)$, then $x$ satisfies the $i_{0}$’th equation, and therefore we must have $x_{j_{0}}=k\in S_{ij_{0}}^{\prime}$. Also, from Lemma B1, $x_{j}\in I_{ij}$, $\forall j\in\mathscr{J}$. Hence, Corollary 3.3 implies that $x$ also satisfies the $i$’th equation. ∎
The Proof of Rule 4.5.
@@
Let $x\in S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)$. From Lemma B1, $x_{j_{0}}\in I_{j_{0}}$ and $x_{j}\in I_{i_{0}j}$, $\forall j\in\mathscr{J}$. But, since $S_{i_{0}j_{0}}^{\prime}=I_{j_{0}}$, then $x_{j_{0}}\in I_{j_{0}}=S_{i_{0}j_{0}}^{\prime}$, and therefore from Corollary 3.3 it follows that $x$ satisfies the $i_{0}$’th equation. ∎
Acknowledgement
The authors wish to express their appreciation for several excellent suggestions for improvements in this paper made by the referees.